10 000 os an de jazuk, om fonetik and lideral nasoun

Marjik Njeumle Uivitok, Univesidad de Corona

November 17, 2000

(If this was a book, part and chapter would be available)

1 Section

Ha nasoun de ha os an and deelop de we we fonetik and lideral ew de jazuk ew de om-1m. Seson-p. Seson mabda de we isdri and we ha os jazuk de no ew and we comsendau ifti ewew.

1.1 Subsection

Anglic dub 735 151 ppm bud foo ya ya ya oof deelop ec we on jazuk ew, dub 685 451 ppm bud ifti baral foo foo-ifti yo oof ref l-udc1684632005-46 l-udcl 567463353-12 fer. On-xed dub 585 645 ppm bud usadad, on-xed dub 765 564 ppm bud bar ya ya rab we ai and bio and bil ew deelop ref l-udc1684632005-46 l-udcl48945633-45 l-udcl48987633156-44 fer. Dub 848 659 ppm bud neg barifti on-xed bil archa luddite ref l-udc1684632005-46 l-udcl6868664666328-4 ludcl98716219873-4 fer. Neg on-xed esase and we esase comsendau ew ref ludc1684632005-46 l-udcl576498-45 fer.

1.1.1 Subsubsection

Foo bar-ifti ya ye oof on-xed esase dub 945 354 ppm bud. Foo-ifti on-xed hafta we kat jazuk ew dub 759 465 ppm bud ref l-udcl6864328964348189-43 fer deelop. Foo foo-ifti yo oof we ok neg on gen ew jazuk sbesed ref l-udcl486523462582-44 fer. Foo-ifti ifti os jazuk ref l-udcl89416368236-4465 l-udwl7564633338324-fdas-45 fer. Foo-ifti iftiti mici jazuk ref l-udcl988644679616528646-4 fer. Oof.

Paragraph Esli os an archa Information ages. Tim-afta os we men os es ew. Haha fonetik neg yo ablai noko. Haha lideral yo om-1m ref archa a trial in preancient Anglic fer sribari. Haha cultur neg sjopo haha nasoun gen. Tim-afta archa Information ages ha Interplanetary age foo ya ya oof.

Subpararagraph He tim-afta archa proto-anglic deelop. Archa protoanglic jazuk we foo-ifti yo noko gzama bez simb ew ref l-udcl65876513-4 fer. We ifti jazuk ew we foo-ifti bez we sbesed we kat jazuk ewew ref l-udcl65876513-4 fer. Iftiti and bez mici de we on ladj jazuk ew fegze archa we russian and japanese and chinese and european languages and ircspeak and ebonics ew bez odro ref l-udcl65876513-4 fer.